# Critical Phase Transition in a Large Language Model

### v1

@@ 3.1 Correlation between Two Tags
$C_{ab}(t,t+\Delta t)=\left\langle\delta_{a,y_{t}}\delta_{b,y_{t+\Delta t}}\right\rangle-\left\langle\delta_{a,y_{t}}\right\rangle\left\langle\delta_{b,y_{t+\Delta t}}\right\rangle,$ (1)
[…]
@@
Figure 2 shows the correlation $C(t,t+\Delta t)$ at $T=0.3$, $1$, and $1.7$ (see App. C for the results at other temperatures). At $T=0.3$, the correlation converges to a positive value as the interval $\Delta t$ increases, whereas the plateau value increases with the position $t$ of the former tag. This is typical behavior when a given system has a long-range order. At the boundary temperature of $T=1$ between low and high temperatures, the correlation decays in a power-law function of $\Delta t$ with a constant prefactor. This is a critical decay with a divergent time scale. Apparently, the correlation at $T=1.7$ also follows a power law. At first sight, this seems to indicate the system is critical at even high temperatures. However, the prefactor for this decay becomes smaller with increasing $t$, meaning that the decay is faster than the critical decay.
To better understand the $T$ dependence of the correlation function, we compute the integrated time correlation:
$\tau_{ab}$ $=\frac{1}{N}\sum_{t,t^{\prime}}C_{ab}(t,t^{\prime})=N\left(\left\langle m_{a}m_{b}\right\rangle-\left\langle m_{a}\right\rangle\left\langle m_{b}\right\rangle\right),$ (2)
where $m_{a}=\sum_{t}\delta_{a,y_{t}}/N$ is the proportion of POS $a$ in a sequence. When the correlation converges to a finite value or follows a critical decay, $\tau_{ab}$ diverges at $N\to\infty$. Therefore, a finite $\tau_{ab}$ in the $N\to\infty$ limit means that the correlation decays qualitatively faster than a critical decay. This quantity allows us to clearly discriminate between the two distinct behaviors from the $N$ dependence of $\tau_{ab}$.
We show the integrated correlation $\tau=\tau_{\text{PROPN},\text{PROPN}}$ as a function of $T$ in Fig. 3 (A). The integrated correlation increases with decreasing temperature, with a peak at $T\approx 1.1$, and then rises again towards lower temperatures. Notably, $\tau$ strongly depends on $N$ in the low-temperature regime, $T\lesssim 1$, while it converges to a finite value at high temperature $T\gtrsim 1.5$ for large $N$, as more clearly seen in $\tau$ as a function of $N$, see Fig. 3 (B). These observations strongly suggest the presence of a critical temperature $T_{c}\approx 1$ below which $\tau$ is infinitely large at $N\to\infty$, meaning that, at low temperatures, the generated sequences are qualitatively different from those at high temperatures. In other words, a phase transition with divergent statistical quantities occurs at the temperature $T_{c}$, as illustrated in Fig. 1 (B).
@@ 3.2 Power spectra of POS sequences
Our numerical results raise the question of what structures emerge in the sequences below $T_{c}$. To answer this question, we compute the power spectrum, that is, the amplitudes of the Fourier modes of the sequences characterized by POS $a$:
$S_{a}(\omega)=N\left(\left\langle\left|\frac{1}{N}\sum_{t}\mathrm{e}^{-2\pi i\omega t}\delta_{a,y_{t}}\right|^{2}\right\rangle-\left|\left\langle\frac{1}{N}\sum_{t}\mathrm{e}^{-2\pi i\omega t}\delta_{a,y_{t}}\right\rangle\right|^{2}\right).$ (3)
The results of $S(\omega)=S_{\text{PROPN}}(\omega)$ at $T=0.3$, $1$, and $1.7$ are shown in Fig. 4 (see App. D for other temperatures). The spectrum at high temperatures is relatively featureless, with a single peak at $\omega=0$. At $T=1$, near $T_{c}$, it has multiple small peaks besides the one at $\omega=0$ for $N=512$, but those peaks do not diverge at $N\to\infty$. With further decreasing temperature, $S(\omega)$ has many peaks that grow with $N$, indicating the presence of long-range order. This spectral behavior clearly demonstrates that the well-known repetitive structures emerge only below the critical temperature. However, the analysis reveals a much richer structure than expected: The sequences indeed have repetitions but with many distinct Fourier modes. This characteristic power spectrum at low temperature is in contrast with simple periodic structures, which have only a few number of dominant Fourier modes.
@@ 3.3 Time evolution of POS sequences
Thus far, we have analyzed the statistical properties of the entire sequences of length $N$ from the beginning. The behavior of correlation $C(t,t+\Delta t)$ has suggested that, on a short time scale, the distribution of the POS tags strongly depends on $t$, and it gradually approaches its stationary state at large $t$ (see Fig. 2). This observation motivates further analysis of the sequences: How the POS tags evolve with time.
We calculate the probability of POS tag at time $t$ being $a$, $v_{a}(t)=\left\langle\delta_{y_{t},a}\right\rangle$. Figure 5 shows $v(t)=v_{\text{PROPN}}(t)$ as a function of time at several temperatures. This not only confirms the aforementioned observation but also reveals that the time scale to reach the stationary value greatly varies with $T$. At temperatures far from $T_{c}$, both higher and lower, $v(t)$ reaches the limiting value rapidly; at low temperatures, it takes a small value, $v\approx 0.1$, while it is a larger one $v\approx 0.4$ at high temperatures. This is in contrast to the fact that the correlation at high temperatures significantly depends on $t$ even at $t=255$, see Fig. 2 (C). Near $T_{c}$, on the other hand, the convergence of $v(t)$ to its stationary state is much slower, indicating critical slowing down. Considering that GPT-2 generates long comprehensible texts around $T_{c}$, we expect that texts are natural only during the transient time before reaching the stationary state and that texts become too repetitive or incomprehensible once $\{v_{a}(t)\}$ reaches the stationary state. If this is indeed the case, then GPT-2 should work well even at lower temperatures if the required sequence length is shorter than the transient time. This view is consistent with the empirical fact that LLMs perform well for practical tasks even when the temperature is lower than $T_{c}\approx 1$ (Renze and Guven, 2024). To reveal the underlying dynamics in the entire 18-dimensional space of $\{v_{a}(t)\}$, we employ principal component analysis (PCA). The result shows that $v_{\text{PROPN}}(t)$ dominates the first principal component (PC), meaning that $v_{\text{PROPN}}(t)$ makes an important contribution in the dynamics of $\{v_{a}(t)\}$. Similar critical behaviors are also observed in the two-dimensional PC space (these results are presented in App. E).
@@ 4 Criticality in Natural Language Corpora
We have shown that generated sequences by GPT-2 have critical properties at $T_{c}\approx 1$. Empirically, GPT-2 generates comprehensible texts around the same point. From these, it is reasonable to consider that natural languages have statistical properties similar to critical GPT-2. Indeed, earlier studies have found power-law decays in correlations across some natural language text and speech corpora (Li, 1989; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023).
@@
To test this idea, we first calculate the perplexity of GPT-2 at various $T$ on OpenWebTextCorpus (OWTC) (Gokaslan and Cohen, 2019), see Fig. 6 (A)  [^3] The perplexity was calculated following the procedure mentioned in the documentation of Transformers (https://huggingface.co/docs/transformers/v4.40.2/en/perplexity).. It becomes minimum at around $T=1$, supporting the above argument. We further compare the correlation function and the integrated correlation of the OWTC with those of the GPT-2 sequences. They all exhibit qualitatively similar behaviors with the GPT-2 sequences near $T_{c}$; a power-law decay of the correlation and the integrated correlation growing with $N$, see Fig. 6 (B) and (C). The time dependence of $v(t)$ for the OWTC is also very similar to that of GPT-2 at $T=0.9$, see Fig. 6(D). Strikingly, both the OWTC and GPT-2 at $T=0.9$ follow almost the same trajectory in this plot over a long time scale, indicating quantitative agreement in their POS distributions. The power spectrum and the dynamics in the two-dimensional PC space also have similar properties (see App. F). Note that the WikiText dataset (Merity et al., 2016) displays statistical properties similar to those of critical GPT-2 as well (see App. G). We thus expect that the criticality observed here is common across different natural language datasets.
5 Discussion and Future Work
@@
In summary, we have studied the effect of the temperature parameter on the GPT-2 sequences and shown that a critical phase transition occurs at $T_{c}\approx 1$. Our results provide a possible picture that, near $T_{c}$, comprehensible sentences emerge only during the transient time regime before reaching the stationary state. We have further demonstrated that the criticality in some statistical quantities appears in natural language corpora as well. We now discuss some implications of our numerical results.
Statistical metrics for evaluating LLMs: If natural languages universally have the critical properties observed in our study, the performance of LLMs might be evaluated using statistical quantities such as the correlations. This simple statistical test needs only multiple sample sequences to see their $N$ dependence, regardless of tasks, corpus, or models. Note that, in practical tasks, it should be sufficient to find a temperature region in which the behavior of sequences is critical within the required sequence length, rather than the exact critical point, as we have argued in Sec. 3.3. Therefore, extensive sampling should not be needed practically. To establish this method, it is important to study the relation between criticality and performance in different tasks.
@@
Critical phenomena and universality class of LLMs and natural languages: We have discussed the critical behaviors in sequences generated by GPT-2 at around $T_{c}$ and shown that GPT-2 and the natural language corpora have common critical behaviors. In the literature of statistical physics, critical phase transitions are classified into universality classes, each of which has a distinct set of critical exponents characterizing the criticality. We expect that the critical transition we have discussed also belongs to a certain universality class, which could be a unique one. Precisely estimating the set of critical exponents for the transition for GPT-2, other LLMs, and multiple natural language corpora is certainly an interesting direction for future work. It is noteworthy that japanese-gpt2-medium (Zhao and Sawada, ; Sawada et al., 2024) have statistical properties qualitatively similar to GPT-2 (see App. H).
@@
Unique statistical properties at high and low temperatures: The statistical behavior in the high-temperature phase has a characteristic behavior that is not common in conventional equilibrium disordered phases: Whereas the integrated correlation is convergent, the decay of correlation is not simply exponential. This behavior is attributed to the decreasing prefactor, which needs to decay sufficiently fast to have a finite $\tau$. Precisely following the $t$ dependence of $C(t,t+\Delta t)$ with fixed $\Delta t$ is needed to address this point.
@@
The low-temperature phase also exhibits an intriguing phenomenon: This phase is definitely ordered in the sense that a long-range time correlation exists. However, unlike conventional ordered phases in equilibrium statistical mechanics, the dynamics of the low-temperature GPT-2 is complex: Sequences have repetitive structures with many peaks in the power spectrum. Up to length $N=512$, the number of peaks grows with $N$, suggesting that the power spectrum would have continuous support of diverging components. If this observation is confirmed, it implies that the low-temperature sequences are chaotic in nature. The temperature dependence of the complex structure is also enigmatic. We have also noticed that, in the $T\to 0$ limit, GPT-2 performs top-$1$ sampling and is deterministic, and the correlation must be strictly zero. Therefore, another phase transition could exist at very low temperature between the repetitive and deterministic phases. We cannot identify this temperature scale because fluctuations are too small to precisely follow the $N$ dependence at this temperature regime. Even more extensive calculations are necessary to study this possible transition. We cannot conclude on the connection between this potential transition and the one argued in Bahamondes (2023).
Approaches to the mechanism of LLMs: As we have already pointed out, the correlation and the power spectrum are peculiar compared to equilibrium statistical-mechanical systems. Whereas we have not fully examined the potential effect of POS tagging at very high temperatures, where generated texts contain many incomprehensible words yet the tagger automatically assigns POS tags, we believe the peculiar phenomena should originate from the distinct features in GPT-2; nonreciprocal and infinite-range interactions implemented with attention mechanisms and highly non-uniform parameters trained on large-scale corpora. Thus, understanding which properties of LLMs cause these phenomena should help reveal the mechanism of LLMs. One possible research direction is an experimental investigation of GPT-2 and other LLMs to identify subnetworks closely related to the phenomena. Another direction is to find a simple mathematical model that reproduces the phenomena observed in this research. In statistical mechanics, mathematical models of machine learning have been studied extensively (Carleo et al., 2019). However, these models do not have critical phase transitions with complex structures at low temperature or apparent power-law decays of correlation at high temperature. This means that existing mathematical models do not capture the properties of LLMs associated with these phenomena.

### v2

@@ 4.1 Correlation between Two Tags
$C_{ab}(t,t+\Delta t)=\mathbb{E}[\delta_{a,y_{t}}\delta_{b,y_{t+\Delta t}}]-\mathbb{E}[\delta_{a,y_{t}}]\mathbb{E}[\delta_{b,y_{t+\Delta t}}],$ (1)
where $a$ and $b$ are POS such as NOUN and VERB, $\delta_{a,b}$ is the Kronecker delta, and $\mathbb{E}[\cdot]$ stands for an average over the POS sequences at a fixed temperature $T$. The dependence of the correlation function on the time interval $\Delta t$ indicates the structure of generated texts. If a text is completely random and disordered, the correlation function decays rapidly with $\Delta t$. In contrast, a text with an order, such as a repetitive structure, exhibits a correlation that remains finite even at large $\Delta t$. Among $18\times 17/2$ pairs of $a$ and $b$, we mainly discuss the case with both $a$ and $b$ being the proper noun (PROPN), which has the largest contribution to the phase transition. This is justified by the fact that other pairs with large contributions show similar behaviors (These points are discussed in App. B in detail). Therefore, we simply refer to $C_{\text{PROPN},\text{PROPN}}$ as $C$. Unless otherwise specified, a similar notation is used for other quantities.
Figure 2 shows the correlation $C(t,t+\Delta t)$ at $T=0.3$, $1$, and $1.7$ (see App. C for the results at other temperatures). At $T=0.3$, the correlation converges to a positive value as the interval $\Delta t$ increases, whereas the plateau value increases with the position $t$ of the former tag. This is typical behavior when a given system has a long-range order. At the boundary temperature of $T=1$ between low and high temperatures, the correlation decays in a power-law function of $\Delta t$, independent of $t$. This is a critical decay with a divergent time scale. Apparently, the correlation at $T=1.7$ also follows a power law. At first sight, this seems to indicate the system is critical at even high temperatures. However, the prefactor for this decay becomes smaller with increasing $t$, meaning that the decay is faster than the critical decay.
@@
To understand the $T$ dependence of the correlation function more systematically, we compute the integrated time correlation:
$\tau_{ab}$ $=\frac{1}{N}\sum_{t,t^{\prime}}C_{ab}(t,t^{\prime})=N\left(\mathbb{E}[m_{a}m_{b}]-\mathbb{E}[m_{a}]\mathbb{E}[m_{b}]\right),$ (2)
where $m_{a}=\sum_{t}\delta_{a,y_{t}}/N$ is the proportion of POS $a$ in a sequence. When the correlation converges to a finite value or follows a critical decay, $\tau_{ab}$ diverges at $N\to\infty$. Therefore, a finite $\tau_{ab}$ in the $N\to\infty$ limit means that the correlation decays qualitatively faster than a critical decay. We can clearly discriminate between the two distinct behaviors from the $N$ dependence of $\tau_{ab}$.
We show the integrated correlation $\tau=\tau_{\text{PROPN},\text{PROPN}}$ as a function of $T$ in Fig. 3 (A). The integrated correlation increases with decreasing temperature, with a peak at $T\approx 1.1$, and then increases again towards lower temperatures. Figure 3 (B) shows the $N$ dependence more clearly. In the low-temperature regime, $T\lesssim 1$, $\tau$ increases algebraically with $N$, suggesting the divergence in the large size limit $N\to\infty$. On the other hand, at high temperatures $T\gtrsim 1$, it saturates to a finite value. These observations strongly suggest the singular behavior of $\tau$: A critical temperature $T_{c}\approx 1$ exists, such that in the large size limit $N\to\infty$, the integrated correlation for $T>T_{c}$ increases as the $T$ approaches $T_{c}$, eventually diverge at $T=T_{c}$. This behavior is singular at $T_{c}$, indicating a phase transition at the temperature. The structure of generated texts should be qualitatively different between the low and high temperature regimes. This behavior is similar to that of the susceptibility in ferromagnetic Ising systems in Fig. 1 (A), except that the integrated correlation diverges even below the critical point. Note that if we consider indices $t$ as spatial positions, sequences $(y_{0},\cdots,y_{N-1})$ can be regarded as a one-dimensional spin configuration of $N$ spins with nonreciprocal and infinite-range interactions. In this interpretation, the integrated correlation defined by Eq. 2 corresponds to the susceptibility.
@@ 4.2 Power spectra of POS sequences
Our numerical results raise the question of what structures emerge in the sequences below $T_{c}$. To answer this question, we compute the power spectrum:
$S_{a}(\omega)=N\left(\mathbb{E}\left[\left|\frac{1}{N}\sum_{t}\mathrm{e}^{-2\pi i\omega t}\delta_{a,y_{t}}\right|^{2}\right]-\left|\mathbb{E}\left[\frac{1}{N}\sum_{t}\mathrm{e}^{-2\pi i\omega t}\delta_{a,y_{t}}\right]\right|^{2}\right).$ (3)
The power spectrum is the amplitudes of the Fourier modes of the sequences characterized by POS $a$. Therefore, its peak at $\omega$ means that a periodic structure with frequency $\omega$ exists.
The results of $S(\omega)=S_{\text{PROPN}}(\omega)$ at $T=0.3$, $1$, and $1.7$ are shown in Fig. 4 (see App. D for other temperatures). The spectrum at high temperatures is relatively featureless, with a single peak at $\omega=0$. At $T=1$, near $T_{c}$, it has multiple small peaks besides the one at $\omega=0$, but those peaks do not diverge at $N\to\infty$. With further decreasing temperature, $S(\omega)$ has many peaks that grow with $N$, indicating the presence of long-range order. This spectral behavior clearly demonstrates that the well-known repetitive structures emerge only below the critical temperature. However, the analysis reveals a much richer structure than expected: The sequences indeed have repetitions but with many distinct Fourier modes. This characteristic power spectrum at low temperature is in contrast with simple periodic structures, which have only a few number of dominant Fourier modes.
@@ 4.3 Time evolution of POS sequences
Thus far, we have analyzed the statistical properties of the entire sequences of length $N$ from the beginning. The behavior of correlation $C(t,t+\Delta t)$ in Fig. 2 has suggested that, at small $t$, the distribution of the POS tags strongly depends on $t$, and it gradually approaches its stationary state at large $t$ as the correlation with the initial time decreases. This observation motivates further analysis of the sequences: How the POS tags evolve with time.
We calculate the probability of POS tag at time $t$ being $a$, $v_{a}(t)=\mathbb{E}[\delta_{y_{t},a}]$. Figure 5 shows $v(t)=v_{\text{PROPN}}(t)$ as a function of time at several temperatures. This not only confirms the aforementioned observation but also reveals that the transient time to reach the stationary value greatly varies with $T$. At temperatures far from $T_{c}$, both higher and lower, $v(t)$ reaches the limiting value rapidly; at low temperatures, it takes a small value, $v\approx 0.1$, while it is a larger one, $v\approx 0.4$, at high temperatures. Near $T_{c}$, on the other hand, the convergence of $v(t)$ to its stationary state is much slower, indicating critical slowing down. Considering that GPT-2 generates long comprehensible texts around $T_{c}$, we expect that texts are natural only during the transient time before reaching the stationary state and that texts become too repetitive or incomprehensible once the POS distribution reaches the stationary state. If this is indeed the case, then GPT-2 should work well even at lower temperatures if the required sequence length is shorter than the transient time. This view is consistent with the empirical fact that LLMs perform well for practical tasks even when the temperature is lower than $T_{c}\approx 1$ (Renze and Guven, 2024).
Since $a$ takes 18 different POS tags, $v_{a}(t)$’s form the 18-dimensional vector ${\bm{v}}(t)$. To reveal the underlying dynamics in the entire 18-dimensional space, we employ principal component analysis (PCA). The result shows that $v_{\text{PROPN}}(t)$ dominates the first principal component (PC), meaning that $v_{\text{PROPN}}(t)$ makes an important contribution in the dynamics of ${\bm{v}}(t)$. In addition, critical behaviors similar to those of $v_{\text{PROPN}}(t)$ are also observed in the two-dimensional PC space. From these results, we can confirm that it is relevant to focus on $v_{\text{PROPN}}(t)$ (The results are presented in App. E).
@@ 5 Criticality in Natural Language Corpora
We have shown that generated sequences by GPT-2 have critical properties at $T_{c}\approx 1$. Empirically, GPT-2 generates comprehensible texts around the same point. From these, it is reasonable to consider that natural languages have statistical properties similar to critical GPT-2. Indeed, earlier studies have found power-law decays in correlations across natural language text and speech corpora (Li, 1989; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023).
@@
To test this idea, we first calculate the perplexity of GPT-2 at various $T$ on OpenWebTextCorpus (OWTC) (Gokaslan and Cohen, 2019) [^4] The perplexity was calculated following the procedure mentioned in the documentation of Transformers (https://huggingface.co/docs/transformers/v4.40.2/en/perplexity).. Figure 6 (A) shows that the perplexity becomes minimum at around $T=1$, meaning that GPT-2 is the closest to the corpus at the point. And then, we compare the statistical properties of the OWTC with those of the GPT-2 sequences. The correlation function and the integrated correlation of the OWTC, shown in Figs. 6 (B) and (C), respectively, exhibit qualitatively similar behaviors with the GPT-2 sequences near $T_{c}$; a power-law decay of the correlation and the integrated correlation growing with $N$. As Fig. 6 (D) presents, the time dependence of $v(t)$ for the OWTC is also very similar to that of GPT-2 at $T=0.9$. Strikingly, both the OWTC and GPT-2 at $T=0.9$ follow almost the same trajectory in this plot over long time, indicating quantitative agreement in their POS distributions. The power spectrum and the dynamics in the two-dimensional PC space also have similar properties (see App. F for the results). Note that the WikiText dataset (Merity et al., 2016) displays statistical properties similar to those of critical GPT-2 as well, as shown in App. G. We thus expect that the criticality observed here is common across different natural language datasets.
6 Discussion
@@
In summary, we have studied the effect of the temperature parameter on the GPT-2 sequences and shown that a critical phase transition occurs at $T_{c}\approx 1$. We have further demonstrated that the criticality in some statistical quantities appears in natural language corpora as well. It is reasonable that LLMs exhibit critical behaviors similar to those of natural languages near the default temperature $T=1$, at which logits are not rescaled. However, the existence of the phase transition is nontrivial. Moreover, the high and low temperature regimes far from $T=1$ show the unexpected behaviors. Our results also provide a possible picture that, near $T_{c}$, comprehensible sentences emerge only during the transient time before reaching the stationary state. We now discuss some implications of our numerical results.
Phase transition in the large context window limit: Since we have analyzed the statistical properties of sequences shorter than the context window of the LLM, the phenomena discussed here should be observed within its scale. Therefore, to increase the system size $N$, the context window must be longer than $N$. The true phase transition in the $N\to\infty$ limit should exist in the limit where the size of models and the length of texts in the training data are infinitely large.
Robustness of the phase transition: The phase transition possibly occurs with varying other parameters, because in crossing a phase boundary in a multi-dimensional parameter space, a phase transition is observed, irrespective of the direction. We also expect that the phase transition exists with other mappings of texts to sequences, other languages, and other LLMs, from the following facts: The power-law decay of correlation is universally observed in different mappings and languages (Li, 1989; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023); The behavior of generated texts with varying temperature is similar with different LLMs, that is, generated texts have repetitions at low temperatures while they are incomprehensible at high temperatures; The effect of details of systems is weak near the critical point due to the divergence of scale. Indeed, in App. H and I, we observe similar phase transitions with a mapping based on characters, and for Japanese GPT-2 (Zhao and Sawada, ; Sawada et al., 2024) and Pythia (Biderman et al., 2023). Note that if the phase transition occurs with different situations, its properties may be different.
@@
Universality class of LLMs and natural languages: We have discussed the critical behaviors in sequences generated by GPT-2 at around $T_{c}$ and shown that GPT-2 and the natural language corpora have common critical behaviors. In the literature of statistical physics, critical phase transitions are classified into universality classes, each of which has a distinct set of critical exponents characterizing the criticality. We expect that the critical phase transition we have discussed also belongs to a certain universality class, which could be a unique one. Precisely estimating the set of critical exponents for the transition for GPT-2, other LLMs, and multiple natural language corpora is certainly an interesting direction for future work.
@@
Unique statistical properties at high and low temperatures: The statistical behavior in the high-temperature phase has a characteristic behavior that is not common in conventional equilibrium disordered phases: Whereas the integrated correlation is convergent, the decay of correlation is not simply exponential. This behavior is attributed to the decreasing prefactor, which needs to decay sufficiently fast to have a finite $\tau$. Precisely following the $t$ dependence of $C(t,t+\Delta t)$ with fixed $\Delta t$ is needed to address this point. Whereas we have not fully examined the potential effect of POS tagging at very high temperatures, where generated texts contain many incomprehensible words yet the tagger automatically assigns POS tags, any such effect should be insignificant at low temperatures and near the critical point, which are our main focus.
@@
The low-temperature phase also exhibits an intriguing phenomenon: This phase is definitely ordered in the sense that a long-range time correlation exists. However, unlike conventional ordered phases in equilibrium statistical mechanics, the dynamics of the low-temperature GPT-2 is complex: Sequences have repetitive structures with many peaks in the power spectrum. Up to length $N=512$, the number of peaks grows with $N$, suggesting that the power spectrum would have continuous support of diverging components. If this observation is confirmed, it implies that the low-temperature sequences are chaotic in nature. We have also noticed that, at $T=0$, GPT-2 performs top-$1$ sampling and is deterministic, and the correlation must be strictly zero. Therefore, another phase transition could exist at very low temperature between the repetitive and deterministic phases. We cannot identify this temperature scale because fluctuations are too small to precisely follow the $N$ dependence at this temperature regime. Even more extensive calculations are necessary to study this possible transition. We cannot conclude on the connection between this potential transition and the one argued in Bahamondes (2023).
These peculiar phenomena should originate from the distinct features in LLMs: nonreciprocal and infinite-range interactions implemented with attention mechanisms and highly non-uniform parameters trained on large-scale corpora. Importantly, although mathematical models of machine learning have been extensively studied in statistical physics (Carleo et al., 2019), existing models do not have critical phase transitions with the unique properties observed in this study.